OpenStudy (anonymous):

Identify the correct categories for the solution to the following equation. I. Natural II. Whole III. Integer IV. Rational V. Irrational x²=36 a.)1 and II b.)III and IV c.)II, III, and IV d.)II and III

OpenStudy (anonymous):

The answer is integer

OpenStudy (anonymous):

X=+-6

OpenStudy (anonymous):

i'm pretty sure it's neither natural nor irrational. it's ineteger because x²=36 -> x=-6 or x=6 so you can eliminate choice a and c

OpenStudy (anonymous):

It can't be Natural, due to the presence of -6

OpenStudy (anonymous):

so the right answer is b

OpenStudy (anonymous):

It can't be whole due to the same reason

OpenStudy (anonymous):

You can call it rational, but integer is a better option

OpenStudy (anonymous):

So the answer having both rational and integer is the right one

OpenStudy (anonymous):

b or c ? hmm . . .

OpenStudy (anonymous):

is it whoLe ?

OpenStudy (anonymous):

No not whole

OpenStudy (anonymous):

Whole means 0,1,2,3......................

OpenStudy (anonymous):

There is no negative in whole

OpenStudy (anonymous):

the answer is b riqht ? cuss its rationaL ?

OpenStudy (anonymous):

yes

OpenStudy (nowhereman):

You can solve that equation in any of those categories, but irrational makes no sense as there are constants (2 and 36) in the equation, that are not irrational. All the others are ok. Of course you will get sometimes more solutions and sometimes less. But the equation itself does not tell you which category you should take. And you could always define a bigger category with still more solutions. (For example you have a solution in ℝ^2: \[\begin{pmatrix}5\\\sqrt{11}\end{pmatrix}^2 =\begin{pmatrix}5\\\sqrt{11}\end{pmatrix}\cdot\begin{pmatrix}5\\\sqrt{11}\end{pmatrix} = 36\] Oh and what your the difference between whole and integer numbers? So why is the answer b and not c?

OpenStudy (nikvist):

Answer is c because \[Z\in Q\]

OpenStudy (nikvist):

Correction \[Z\subset Q\]

OpenStudy (anonymous):

yes i think you're right i googled about whole number and it's almost the same as integer so the right answer is c

OpenStudy (anonymous):

okay thanks ! :D

OpenStudy (anonymous):

Could you provide me some help on this Well, here is my next question http://openstudy.com/updates/4d9abb520b248b0b471ce9e2?source=email#/updates/4d9ac0cf0b248b0bdc1fe9e2

OpenStudy (anonymous):

Is this the question?

OpenStudy (anonymous):

I am sorry,

OpenStudy (anonymous):

This not the question

OpenStudy (anonymous):

There is some problem with the site

OpenStudy (anonymous):

I am providing the link.

OpenStudy (anonymous):

Yes this is the link